-15x+y=55
find y and x intercept

Answer by sudhanshu_kmr(1152)


-15x+y=55
put y = 0 to find x intercept
-15x = 55
=> x = -11/3

now, put x = 0 to find y intercept..
y= 55

x intercept = -11/3, y-intercept = 55